Exemple :- o given R, lokas, R2 = 12.5kan, C = IMF, (2= 2uF. Determine impulse responte hlt) of siegure. - . Vies)

Question is given inside the pic

May I have a detailed explanation of the problem?

and how do we get the equivalent resistance for R1 II 1/C1S and the 1 + 1/C2S

PLEASE A DETAILED EXPLANATION
Answer #1

2 2 Re=12.5kn. =10Kn 12.5 kv. Pris 245 245 - Ve(s) 5 = 1 4F C2 =24f. → Z(s) = R XT AS a R 1 +RICS 2115) To 1.09 VG) 2017 YalcH(s) = 1 + (-2x10²s) 2.5xits +3.55 Xolta H(S) = 1 2005 S 2.55 + 5505 +10000 -550 -4x2.5 X 10000 2x2.5 S, = -200, S2 = -20 =)